# DifFlow3D: Toward Robust Uncertainty-Aware Scene Flow Estimation with Iterative Diffusion-Based Refinement

### v2

@@ Abstract
Scene flow estimation, which aims to predict per-point 3D displacements of dynamic scenes, is a fundamental task in the computer vision field. However, previous works commonly suffer from unreliable correlation caused by locally constrained searching ranges, and struggle with accumulated inaccuracy arising from the coarse-to-fine structure. To alleviate these problems, we propose a novel uncertainty-aware scene flow estimation network (DifFlow3D) with the diffusion probabilistic model. Iterative diffusion-based refinement is designed to enhance the correlation robustness and resilience to challenging cases, e.g., dynamics, noisy inputs, repetitive patterns, etc. To restrain the generation diversity, three key flow-related features are leveraged as conditions in our diffusion model. Furthermore, we also develop an uncertainty estimation module within diffusion to evaluate the reliability of estimated scene flow. Our DifFlow3D achieves state-of-the-art performance, with $6.7\%$ and $19.1\%$ EPE3D reduction respectively on FlyingThings3D and KITTI 2015 datasets. Notably, our method achieves an unprecedented millimeter-level accuracy ($0.0089$m in EPE3D) on the KITTI dataset. Additionally, our diffusion-based refinement paradigm can be readily integrated as a plug-and-play module into existing scene flow networks, significantly increasing their estimation accuracy. Codes will be released later.
@@ 1 Introduction
As a fundamental task in computer vision, scene flow refers to the 3D motion field estimated from consecutive images or point clouds. It provides the low-level perception of dynamic scenes and has various down-stream applications, such as autonomous driving [21], pose estimation [9], and motion segmentation [1]. Early works focus on employing stereo [12] or RGB-D images [10] as input. With the increasing application of 3D sensors, eg. LiDAR, recent works commonly take point clouds directly as input.
@@
As a pioneering work, FlowNet3D [16] extracts hierarchical features with PointNet++ [25], and then regresses the scene flow iteratively. PointPWC [42] further improves it with the Pyramid, Warping, and Cost volume structure [31]. HALFlow [35] follows them and introduces the attention mechanism for better flow embedding. However, these regression-based works commonly suffer from unreliable correlation and local optimum problems [17]. Reasons are mainly two folds: (1) K Nearest Neighbor (KNN) is utilized for searching point correspondences in their networks, which can not take into account correct yet distant point pairs and also has matching noise [7]. (2) Another potential problem arises from the coarse-to-fine structure widely used in previous works [16, 42, 35, 36]. Basically, an initial flow is estimated in the coarsest layer and then iteratively refined in higher resolutions. However, the performance of flow refinement highly relies on the reliability of initial coarse flow, since subsequent refinement is typically constrained within small spatial ranges around the initialization.
@@
To address the unreliability issue, 3DFlow [36] designs an all-to-all point-gathering module with backward validation. Similarly, Bi-PointFlowNet [4] and its extension MSBRN [5] propose a bi-directional network with the forward-backward correlation. IHNet [38] leverages a recurrent network with high-resolution guidance and a re-sampling scheme. However, these networks mostly struggle with large computational costs due to their bi-directional association or recurrent iteration. In this paper, we discover that diffusion models can also enhance the correlation reliability and resilience to matching noise due to its denoising intrinsic (Fig. 1). Motivated by the findings in [30] that injecting random noise is helpful to jump out of local optimum, we re-formulate the deterministic flow regression task with a probabilistic diffusion model as illustrated in Fig. 2. Moreover, our method can serve as a plug-and-play module in previous scene flow networks, which is more general and almost has no computation cost increase (Section 4.5).
@@
However, it is rather challenging to leverage a generative model in our task, due to the inherent generation diversity of diffusion models. Unlike the point cloud generation task requiring diverse output samples, scene flow prediction is a deterministic task that calculates precise per-point motion vectors. To tackle this problem, we leverage strong condition information to restrict the diversity and effectively control the generated flow. Specifically, a coarse sparse scene flow is first initialized, and then the flow residuals are iteratively generated with diffusion. In each diffusion-based refinement layer, we utilize coarse flow embedding, cost volume, and geometry encoding as conditions. In this case, diffusion is applied to actually learn a probabilistic mapping from conditional inputs to the flow residual.
@@ 2 Related Work
Scene Flow Estimation. As the counterpart of optical flow in 3D vision, scene flow has witnessed remarkable advances in recent years. FlowNet3D [16] is a pioneering work which firstly extracts point features with PointNet++ [25], and then regresses the per-point motion vector hierarchically. PointPWC [42] further improves it with the Pyramid, Warping, and Cost volume structure [31] in a patch-to-patch manner. FLOT [24] proposes an optimal transport module to establish the cross-frame association. PV-RAFT [40] fuses the point and voxel representations for both local and global feature extraction. To tackle the unreliable correlation, 3DFlow, Bi-PointFlowNet, and MSBRN [36, 4, 5] design the bi-directional correlation layer. IHNet [38] develops an iterative hierarchical network guided by estimated high-resolution information to tackle the mismatches. PT-FlowNet [7] proposes a transformer-based pipeline to capture long-range dependencies. In this paper, we rethink scene flow estimation with a different insight: How to find a more universal method improving the robustness and reliability of previous regression-based methods? On the one hand, we introduce a denoising diffusion model with strong conditions. On the other hand, uncertainty estimation is proposed to evaluate the reliability of dense flow matching.
Diffusion models in 3D vision. Recently, diffusion model [28, 11] has gained significant attention, which gradually removes noise from a Gaussian distribution to the sample data distribution by learning from its reverse process. It has yielded great advancements in image generation [27, 26], video synthesis [6, 43], and static point cloud generation and completion [18, 34]. DPM [18] first applies the diffusion model conditioned on a shape latent to point cloud generation. PVD [44] proposes a 3D shape generation and completion network through point-voxel diffusion. LION [34] combines VAE and two latent diffusion models (feature-based and point-based) to further enhance the generation abilities. However, it remains unclear whether diffusion can model dynamic point clouds, i.e., scene flow.
Uncertainty estimation. Uncertainty estimation has been widely explored in stereo matching, Multi-View Stereo (MVS), and optical flow fields. UCS-Net [3] proposes an uncertainty-aware cascaded network where uncertainty interval is inferred from depth probabilities to progressively sub-divide local depth ranges. SEDNet [2] designs a loss function for joint disparity and uncertainty estimation in deep stereo matching. ProbFlow [39] proposes to jointly estimate the optical flow and its uncertainty. Similarly, PDCNet [33] enhances ProbFlow by a new training strategy in terms of self-supervised training. Basically, the uncertainty of scene flow estimation has been rarely studied before. In this paper, we design a per-point uncertainty to evaluate the reliability of our estimated scene flow.
@@ 3.2 Diffusion-based Scene Flow Refinement
After obtaining the initial scene flow in the coarsest layer, our network will refine the denser scene flow using diffusion model from coarse to fine. It is non-trivial to directly recover scene flow from the pure noise due to the significant difference between their data distributions. Since coarse scene flow has been already obtained in the above flow initialization, what we exactly need is the flow residual between initialized scene flow and the ground truth. Therefore, we formulate the scene flow residual as the diffusion latent variable, and multi-scale flow residuals are generated from the reverse process of diffusion model iteratively.
@@
However, there still remain challenges since scene flow estimation requires the precise value and direction, rather than diverse outputs as in the point cloud generation task. Therefore, we constrain the generation diversity and control the reverse process by the powerful condition information $C$. In this way, the training objective of flow residual can be represented by:
@@ 3.3 Uncertainty Estimation
where ${u}_{0}^{l}$ indicates the constructed uncertainty ground truth. $E_{1}$ and $E_{2}$ decay with the convergence of network, which constrain the uncertainty intervals by absolute and relative estimation errors. As in Fig. 4, uncertainty intervals represent the reliability range of each estimated flow and progressively encourage the estimated flow to optimize toward the ground truth. The supervision of uncertainty residual is similar to flow residual in the diffusion:
@@ 3.4 Overall Architecture of DifFlow3D
We adopt the coarse-to-fine structure to generate the scene flow residuals and uncertainty residuals at different scales. The inputs of our diffusion-based refinement at layer $l$ are coarse sparse scene flow $sf^{l}$, coarse sparse uncertainty $u^{l}$ and condition signals $C^{l}$. Firstly, coarse sparse scene flow $sf^{l}$ is up-sampled to produce the coarse dense scene flow $sf^{dense}$. Similarly, coarse sparse uncertainty $u^{l}$ is up-sampled as coarse dense uncertainty $u^{dense}$. Dense flow residual $\hat{s}_{0}^{l}$, uncertainty residual $\hat{u}_{0}^{l}$ are then refined from the denoising network $M_{\theta}$. The refined dense scene flow (or uncertainty) at layer $l$ are generated by the per-point addition between coarse dense flow (or uncertainty) and its corresponding residual. Finally, the refined dense flow and uncertainty will be the inputs of layer $l+1$:
@@ 3.5 Training Objective
Ours 0.0140 0.9776 0.9933 0.0479 0.7614 0.9759 0.0089 0.9813 0.9930 0.0825 0.3305 0.9925
Table 1: Comparison results on FlyingThings3D and KITTI datasets without occlusion [8]. The best results are in bold. Our method has $\bf{6.7}\%$ and $\bf{19.1\%}$ EPE3D reduction respectively compared with previous works.
@@ 4.1 Datasets and Implementation Details
We follow previous works [16, 8, 24, 40, 36, 4] to train our DifFlow3D on synthetic FlyingThings3D dataset [19], and evaluate the model on both FlyingThings3D and KITTI [20] datasets. All the experiments are conducted on a single RTX 3090 GPU. The Adam optimizer is adopted with $\beta_{1}=0.9,\beta_{2}=0.999$. Initial learning rate is set as 0.001 and decays exponentially with the rate $\gamma=0.5$. We leverage DDIM [29] with the sampling step = 1, and $T$ = 1000. Refinement layer number $K$ is 3. Gaussian noise is set as $\sigma$ = 1. $E_{1}$ and $E_{2}$ are initially set as 0.5, and decay until 0.02. Our evaluation metrics include EPE3D (m), Acc3DS, Acc3DR, Outliers, EPE2D (px), and Acc2D, commonly used in previous works [16, 36, 4]. More implementation details are provided in the supplementary materials.
@@
Ours Full 0.029 0.959 0.975 0.108
Table 2: Comparison results on FlyingThings3D and KITTI datasets with occlusion [16]. “Self” and “Full” respectively mean self-supervised and fully-supervised training. Our method outperforms previous works by $\bf{11.3}\%$ and $\bf{34.1}\%$ in terms of EPE3D.
@@ 4.2 Comparison with State-of-the-Art Methods
To comprehensively demonstrate the superiority of our method, we show comparison results with two pre-process approaches: without occlusion [8] and with occlusion [16].
Comparison on point clouds without occlusion. We compare our DifFlow3D with a series of state-of-the-art (SOTA) methods on both FlyingThings3D [19] and KITTI scene flow [20] datasets without occlusion. Table 1 demonstrates that our DifFlow3D outperforms all previous methods in terms of both 3D and 2D metrics on two datasets. Compared with the recent SOTA method MSBRN [5], our model reduces the End-point-error (EPE3D) by 6.7% and 19.1% respectively on FlyingThings3D and KITTI datasets. It is worth noting that DifFlow3D also has the best generalization capability. Only trained on Flyingthings3D synthetic dataset, our method firstly achieves millimeter-level (0.0089m) EPE3D metric on the KITTI dataset.
[…]
@@
3DFlow+DSFR Quarter 0.0297 (normal-$\downarrow$ 6.3%) 0.9207 0.9785 0.1548 1.6344 0.9188 0.0316 (normal-$\downarrow$ 4.8%) 0.9028 0.9634 0.1604 1.2247 0.9413
@@
3DFlow+DSFR Complete 0.0242 (normal-$\downarrow$13.9%) 0.9494 0.9860 0.1166 1.3201 0.9459 0.0251 (normal-$\downarrow$18.8%) 0.9398 0.9793 0.1302 0.9761 0.9686
@@
BPF+DSFR Complete 0.0252 (normal-$\downarrow$10.0%) 0.9388 0.9824 0.1251 1.5057 0.9390 0.0230 (normal-$\downarrow$23.3%) 0.9467 0.9816 0.1229 0.9647 0.9747
@@
MSBRN+DSFR Complete 0.0140 (normal-$\downarrow$ 6.7%) 0.9776 0.9933 0.0479 0.7614 0.9759 0.0089 (normal-$\downarrow$19.1%) 0.9813 0.9930 0.0825 0.3305 0.9925

### v3

@@ Abstract
Scene flow estimation, which aims to predict per-point 3D displacements of dynamic scenes, is a fundamental task in the computer vision field. However, previous works commonly suffer from unreliable correlation caused by locally constrained searching ranges, and struggle with accumulated inaccuracy arising from the coarse-to-fine structure. To alleviate these problems, we propose a novel uncertainty-aware scene flow estimation network (DifFlow3D) with the diffusion probabilistic model. Iterative diffusion-based refinement is designed to enhance the correlation robustness and resilience to challenging cases, e.g. dynamics, noisy inputs, repetitive patterns, etc. To restrain the generation diversity, three key flow-related features are leveraged as conditions in our diffusion model. Furthermore, we also develop an uncertainty estimation module within diffusion to evaluate the reliability of estimated scene flow. Our DifFlow3D achieves state-of-the-art performance, with $24.0\%$ and $29.1\%$ EPE3D reduction respectively on FlyingThings3D and KITTI 2015 datasets. Notably, our method achieves an unprecedented millimeter-level accuracy ($0.0078$m in EPE3D) on the KITTI dataset. Additionally, our diffusion-based refinement paradigm can be readily integrated as a plug-and-play module into existing scene flow networks, significantly increasing their estimation accuracy. Codes are released at https://github.com/IRMVLab/DifFlow3D.
@@ 1 Introduction
As a fundamental task in computer vision, scene flow refers to the 3D motion field estimated from consecutive images or point clouds. It provides the low-level perception of dynamic scenes and has various down-stream applications, such as autonomous driving [27, 10], SLAM [52, 55, 20], and motion segmentation [1, 6]. Early works focus on employing stereo [13] or RGB-D images [11] as input. With the increasing application of 3D sensors, e.g. LiDAR, recent works commonly take point clouds directly as input.
@@
As a pioneering work, FlowNet3D [22] extracts hierarchical features with PointNet++ [32], and then regresses the scene flow iteratively. PointPWC [51] further improves it with the Pyramid, Warping, and Cost volume structure [39]. HALFlow [43] follows them and introduces the attention mechanism for better flow embedding. However, these regression-based works commonly suffer from unreliable correlation [21] and local optimum problems [23]. Reasons are mainly two folds: (1) K Nearest Neighbor (KNN) is utilized for searching point correspondences in their networks, which can not take into account correct yet distant point pairs and also has matching noise [8]. (2) Another potential problem arises from the coarse-to-fine structure widely used in previous works [22, 51, 43, 44]. Basically, an initial flow is estimated in the coarsest layer and then iteratively refined in higher resolutions. However, the performance of flow refinement highly relies on the reliability of initial coarse flow, since subsequent refinement is typically constrained within small spatial ranges around the initialization.
@@
To address the unreliability issue, 3DFlow [44] designs an all-to-all point-gathering module with backward validation. Similarly, BPF [4] and its extension MSBRN [5] propose a bi-directional network with the forward-backward correlation. IHNet [47] leverages a recurrent network with high-resolution guidance and a re-sampling scheme. However, these networks mostly struggle with large computational costs due to their bi-directional association or recurrent iteration. In this paper, we discover that diffusion models can also enhance the correlation reliability and resilience to matching noise due to its denoising intrinsic (Fig. 1). Motivated by the findings in [38] that injecting random noise is helpful to jump out of local optimum, we re-formulate the deterministic flow regression task with a probabilistic diffusion model as illustrated in Fig. 2. Moreover, our method can serve as a plug-and-play module in previous scene flow networks, which is more generic and almost has no computational overhead (Section 4.5).
@@
However, it is rather challenging to leverage a generative model in our task, due to the inherent generation diversity of diffusion models. Unlike the point cloud generation task requiring diverse output samples, scene flow prediction is a deterministic task that calculates precise per-point motion vectors. To tackle this problem, we leverage strong condition information to restrict the diversity and effectively control the generated flow. Specifically, a coarse sparse scene flow is first initialized, and then the flow residuals are iteratively generated with diffusion. In each diffusion-based refinement layer, coarse flow embedding, cost volume, and geometry encoding are all utilized as conditions. In this case, diffusion is applied to learn a probabilistic mapping from conditional inputs to the flow residual.
@@ 2 Related Work
Scene Flow Estimation. As the counterpart of optical flow in 3D vision, scene flow has witnessed remarkable advances in recent years. FlowNet3D [22] is a pioneering work that firstly extracts point features with PointNet++ [32], and then regresses the per-point motion vector hierarchically. PointPWC [51] further improves it with the Pyramid, Warping, and Cost volume structure [39] in a patch-to-patch manner. FLOT [31] proposes an optimal transport module to establish the cross-frame association. PV-RAFT [49] fuses the point and voxel representations for both local and global feature extraction. To tackle the unreliable correlation, 3DFlow, Bi-PointFlowNet, and MSBRN [44, 4, 5] design the bi-directional correlation layer. SFGAN [45] proposes a scene flow network with the generative adversarial network (GAN). DELFlow [30] proposes a dense efficient flow estimation network for large-scale points. PT-FlowNet [8] proposes a transformer-based pipeline to capture long-range dependencies. There are also some self-supervised methods [1, 19, 14, 35, 17, 15], without requiring ground truth flow annotations. In this paper, we rethink scene flow estimation with a different insight: How to find a more generic method improving the robustness and reliability of previous regression-based methods? On the one hand, we introduce a denoising diffusion model with strong conditions. On the other hand, uncertainty estimation is proposed to evaluate the reliability of dense flow matching.
Diffusion models in 3D vision. Recently, diffusion model [36, 12] has gained significant attention, which gradually removes noise from a Gaussian distribution to the sample data distribution by learning from its reverse process. It has yielded great advancements in image generation [34, 33], video synthesis [7, 53], and static point cloud generation and completion [24, 42]. DPM [24] first applies the diffusion model conditioned on a shape latent to point cloud generation. PVD [54] proposes a 3D shape generation and completion network through point-voxel diffusion. LION [42] combines VAE and two latent diffusion models (feature-based and point-based) to further enhance the generation abilities. However, it remains unclear whether diffusion can model dynamic point clouds, i.e. scene flow.
Uncertainty estimation. Uncertainty estimation has been widely explored in stereo matching, Multi-View Stereo (MVS), and optical flow fields. UCS-Net [3] proposes an uncertainty-aware cascaded network where uncertainty interval is inferred from depth probabilities to progressively sub-divide local depth ranges. SEDNet [2] designs a loss function for joint disparity and uncertainty estimation in deep stereo matching. ProbFlow [48] proposes to jointly estimate the optical flow and its uncertainty. Similarly, PDCNet [41] enhances ProbFlow with a new training strategy in terms of self-supervised training. Nevertheless, the uncertainty of scene flow estimation has been rarely studied before. In this paper, we design a per-point uncertainty to evaluate the reliability of our estimated flow.
@@ 3.2 Diffusion-based Scene Flow Refinement
It is non-trivial to directly recover scene flow from pure noise due to the significant discrepancies between their data distributions. Since coarse scene flow has been already obtained in the above flow initialization, what we exactly need is the flow residual between the initialized scene flow and the ground truth. Therefore, we formulate the scene flow residual as the diffusion latent variable, and multi-scale flow residuals are generated from the reverse process of diffusion model iteratively.
@@
However, there still remain challenges since scene flow estimation requires precise values and directions, rather than diverse outputs as in the point cloud generation task. Therefore, we constrain the generation diversity and control the reverse process by the powerful condition information $C$. In this way, the training objective of flow residual can be represented by:
@@ 3.3 Uncertainty Estimation
where ${u}_{0}^{l}$ indicates the constructed uncertainty ground truth. $E_{1}$ and $E_{2}$ decay with the convergence of the network, which constrain the uncertainty intervals by absolute and relative estimation errors. As in Fig.4, uncertainty intervals represent the reliability range of each estimated flow and progressively encourage the estimated flow to optimize toward the ground truth. The supervision of uncertainty residual is similar to flow residual in the diffusion:
@@ 3.4 Overall Architecture of DifFlow3D
We adopt the coarse-to-fine structure to generate the scene flow residuals and uncertainty residuals at different scales. The inputs of our diffusion-based refinement at layer $l$ are coarse sparse scene flow $sf^{l}$, coarse sparse uncertainty $u^{l}$ and condition signals $C^{l}$. Firstly, coarse sparse scene flow $sf^{l}$ is up-sampled to produce the coarse dense scene flow $sf^{dense}$. Similarly, coarse sparse uncertainty $u^{l}$ is up-sampled as coarse dense uncertainty $u^{dense}$. Dense flow residual $\hat{s}_{0}^{l}$, uncertainty residual $\hat{u}_{0}^{l}$ are then refined from the denoising network $M_{\theta}$. The refined dense scene flow (or uncertainty) at layer $l$ is generated by the per-point addition between coarse dense flow (or uncertainty) and its corresponding residual. Finally, the refined dense flow and uncertainty will be the inputs of layer $l+1$:
@@ 3.5 Training Objective
Ours 0.0114 0.9836 0.9949 0.0350 0.6220 0.9824 0.0078 0.9817 0.9924 0.0795 0.2987 0.9932
Table 1: Comparison results on FlyingThings3D and KITTI datasets without occlusion [9]. The best results are in bold. Our method has $\bf{24.0}\%$ and $\bf{29.1\%}$ EPE3D reduction respectively compared with previous works.
@@ 4.1 Datasets and Implementation Details
We follow previous works [22, 9, 31, 49, 44, 4] to train our DifFlow3D on synthetic FlyingThings3D dataset [25], and evaluate the model on both FlyingThings3D and KITTI [26] datasets. All the experiments are conducted on a single RTX 3090 GPU. The Adam optimizer is adopted with $\beta_{1}=0.9,\beta_{2}=0.999$. The learning rate is initially set as 0.0001 and reduced by 0.8 every 10 epochs. We leverage DDIM [37] with the total timestamp $T$ = 1000. Refinement layer number $K$ is 3. Gaussian noise is set as $\sigma$ = 1. We imitate the learning rate step decay strategy for $E_{1}$ and $E_{2}$, which are initially set as 0.5, and decay with the rate 0.8 until 0.02. Our evaluation metrics include EPE3D (m), Acc3DS, Acc3DR, Outliers, EPE2D (px), and Acc2D, commonly used in previous works [22, 44, 4].
@@
Ours Full 0.031 0.955 0.966 0.108
Table 2: Comparison results on FlyingThings3D and KITTI datasets with occlusion [22]. “Self" and “Full" respectively mean self-supervised and fully-supervised training. Our method outperforms previous works by $\bf{18.9}\%$ and $\bf{29.5}\%$ in terms of EPE3D.
@@ 4.2 Comparison with State-of-the-Art Methods
To comprehensively demonstrate the superiority of our method, we show comparison results with two pre-process settings: without occlusion [9] and with occlusion [22].
Comparison on point clouds without occlusion. We compare our DifFlow3D with a series of state-of-the-art (SOTA) methods on both FlyingThings3D [25] and KITTI scene flow [26] datasets without occlusion. Table 1 demonstrates that our DifFlow3D outperforms all previous methods in terms of both 3D and 2D metrics on two datasets. Compared with the recent SOTA method MSBRN [5], our model reduces the End-point-error (EPE3D) by 24.0% and 29.1% respectively on FlyingThings3D and KITTI datasets. It is worth noting that DifFlow3D also has the best generalization capability. Only trained on the Flyingthings3D synthetic dataset, our method firstly achieves millimeter-level (0.0078m) EPE3D metric on the KITTI dataset.
Comparison on point clouds with occlusion. We also evaluate our model on the datasets with occlusion prepared by [22]. Experiment results in Table 2 show that our DifFlow3D still surpasses all previous methods on both two datasets. Notably, our evaluation results on the KITTI dataset outperform previous methods by a large margin. Only trained on FT3DO, our method reduces EPE3D by $29.5\%$ and outliers by $48.1\%$ on KITTI. This demonstrates the excellent generalization capability of our method.
@@
3DFlow+DSFR Quarter 0.0297 ($\downarrow$ 6.3%) 0.9207 0.9785 0.1548 1.6344 0.9188 0.0316 ($\downarrow$ 4.8%) 0.9028 0.9634 0.1604 1.2247 0.9413
@@
3DFlow+DSFR Complete 0.0242 ($\downarrow$13.9%) 0.9494 0.9860 0.1166 1.3201 0.9459 0.0251 ($\downarrow$18.8%) 0.9398 0.9793 0.1302 0.9761 0.9686
@@
BPF+DSFR Complete 0.0247 ($\downarrow$11.8%) 0.9390 0.9840 0.1192 1.3749 0.9468 0.0265 ($\downarrow$11.7%) 0.9355 0.9672 0.1290 1.0527 0.9637
@@
MSBRN+DSFR Complete 0.0114 ($\downarrow$ 24.0%) 0.9836 0.9949 0.0350 0.6220 0.9824 0.0078 ($\downarrow$29.1%) 0.9817 0.9924 0.0795 0.2987 0.9932